For retirement, you decide to deposit $8805 at the end of each year and you will increase your deposit by $214 per year. How much will you have at the end of 20 years if the bank pays 2.5% compounded annually? Enter your answer as follows: 12345 Round your answer. Do not use a dollar sign ($), any commas() or a decimal point)
Answer #1

Annual deposit = $8805

increase in deposit = $214

i = 2.5%

n = 20 years

we will find present worth first and then the future worth for easier calculations.

pw = annual deposit(p/a,i,n) + increase in deposit(p/g,i,n)

pw = 8805(p/a,2.5%,20) + 214(p/g,2.5%,20)

pw = 8805 * 15.5892 + 214 * 135.3497

pw = 137,262.906 + 28,964.8358

pw = 166,227.7418

fw = pw(f/p,i,n)

fw = 166,227.7418 (f/p,2.5%,20)

fw = 166,227.7418 * 1.6386

fw = $ 253,883.8335716

so at the end of 20 years he will have $253,883.8335716